Ceelo Company purchased (at a cost of $12,510) and used 1,800 pounds of materials during May. Ceelo’s standard cost of materials per unit produced is based on 1 pounds per unit at a cost $7.00 per pound. Production in May was 1,720 units.

(a)

Compute the total, price, and quantity variances for materials. (Round answers to 0 decimal places, e.g. 125.)

Total Materials Variance $ UnfavorableFavorableNeither favorable nor unfavorable
Materials Price Variance $ UnfavorableFavorableNeither favorable nor unfavorable
Materials Quantity Variance $ UnfavorableFavorableNeither favorable nor unfavorable
Answer #1

Total material variance = (1720*1*7)-12510 = 470 Unfavorable

Material price variance = (7*1800-12510) = 90 Favorable

Material quantity variance = (1720-1800)*7 = 560 Unfavorable